In a darts competition, Susan scored 147 more than John and three times as much as Elizabeth. If their combined score came to 336 find their individual scores.
Thanks in advance!

Answer by AnlytcPhil(1810)

In a darts competition, Susan scored 147 more than John and three times as
much as Elizabeth. If their combined score came to 336 find their individual
scores.

>>...Susan scored 147 more than John...<<

S = J + 147

>>...Susan scored...three times as much as Elizabeth...<<

S = 3E

>>...their combined score came to 336...<<

S + J + E = 336

This gives us this system of equations:


S = J + 147
S = 3E
S + J + E = 336

Can you solve that system?  If not re-post.

Answers are S = 207, J = 60, E = 69

Edwin
AnlytcPhil@aol.com